OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -13, -9, -5, ... hellppppp pleaseee

sam (.sam.):

First thing you have to know is whether it is A.P or G.P

OpenStudy (anonymous):

a.p or g.p?

sam (.sam.):

Arithmetic progression Geometric progression

OpenStudy (anonymous):

Arithmetic

sam (.sam.):

From here you can tell that it's AP, yes

OpenStudy (anonymous):

find the difference between each number (i.e. 4), set up initial progression 4n+b then work out what b is by calculation -17 = 4n+b b=-21 so your progression is 4n-21

sam (.sam.):

the common difference is d=-13-(-17)=4

OpenStudy (anonymous):

ohhh okay thank you. could you help me with another one?

OpenStudy (anonymous):

yeah

sam (.sam.):

Then follow what @epicfail said

OpenStudy (anonymous):

so I got 17 + 4(n - 1)

OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 2 - 6 + 18 - 54 + ...